OpenStudy (anonymous):

The equation of a line is y = 2x + 3. What is the equation of the line that is parallel to the first line and passes through (2, –1)? A. 4x – 2y = –6 B. y = 2x – 5 C. y = 3x + 4 D. 2x + y = –1

OpenStudy (ankitshaw):

parallel lines means slope same so the slope of the reqired line will also be 2. y- (-1) = 2(x-2) is the req line (y- y coordinate) = slope(x-x coordinate) y+1=2x-4 ..... y= 2x - 5

OpenStudy (anonymous):

ankitshaw, I appreciate it! :)